OpenStudy (anonymous):

At the market, 777 apples cost \$4.55$4.55dollar sign, 4, point, 55. How much do 999 apples cost?

OpenStudy (texaschic101):

If I am understanding this correctly......777 apples cost $ 4.55....and you want to know how much 999 apples cost... I am going to set up a proportion.. 777/4.55 = 999/x --- (777 to 4.55 = 999 to x) cross multiply 777x = (4.55)(999) 777x = 4545.45 x = 4545.45/777 x = 5.85 so 999 apples cost $ 5.85